# Exciton-Polariton Dynamics in Multilayered Materials

### v1

@@ Abstract
Coupling excitons with quantized radiation has been shown to enable coherent ballistic transport at room temperature inside optical cavities. Previous theoretical works employ a simple description of the material, depicting it as a single layer placed in the middle of the optical cavity, thereby ignoring the spatial variation of the radiation in the cavity quantization direction. In contrast, in most experiments, the optical cavity is filled with organic molecules or multilayered materials. Here, we develop an efficient mixed-quantum-classical approach, introducing a bright layer description, to simulate the exciton-polariton quantum dynamics. Our simulations reveal that, for the same Rabi splitting, a multilayer material extends the quantum coherence lifetime and enhances transport compared to a single-layer material. We find that this enhanced coherence can be traced to a synchronization of phonon fluctuations over multiple layers, wherein the collective light-matter coupling in a multilayered material effectively suppresses the phonon-induced dynamical disorder.
@@
However, these theoretical works adopt a simplified model system that describes the material as a single layer placed at the center of the optical cavity (see Fig. 1a). This is in contrast to most of the experiments [2, 3, 6, 5, 15, 16] where the space between the reflective mirrors is filled with molecules or multiple layers of materials, as schematically illustrated in Fig. 1b. Consequently, the density of states in these two setups (i.e., multilayered vs. single-layered) are drastically different (see insets in Fig. 1c-d), with the multilayered material featuring an ensemble of optically dark bands that are absent in a single-layered material when coupled to cavity. Therefore, it is anticipated that the polariton quantum dynamics in these two setups will differ. However, the role of a multilayer configuration in polaritonic quantum dynamics has remained unexplored, which is the focus of this present work.
@@
Figure 1: Schematics and band structures of single-layer vs. multilayer materials. Illustration of (a) a single layer material and (b) multilayered material in an optical cavity, respectively. The exciton-polariton band structures of (c) single-layer and (d) multilayer materials in the absence of phonons ($\gamma=0$). In a multilayer setup, in addition to the polariton bands, an ensemble of dark material bands exist. These dark bands increase the density of states (DOS) near the onsite energy value, which are plotted as insets within (c)-(d).
@@
Simulating exciton-polariton dynamics in multilayered material is an extremely challenging task, as the number of required unit cells in a single layer exceeds $\sim 10^{4}$, thus requiring us to perform a quantum dynamics simulation involving $\sim 10^{6}$ states for $\sim 10^{2}$ layers, typical in these systems. To resolve this fundamental challenge, here we introduce an efficient mixed quantum-classical approach that uses a bright-layer description of the light-matter interaction and propagates the dynamics utilizing a split-operator approach. Using this convenient theoretical tool, we investigate the polaritonic quantum dynamics in multilayered materials. We discover a phonon fluctuation synchronization effect where the exciton-polaritons experience much lower phonon fluctuations and remain coherent for a longer time. We show that this effect originates from the delocalized nature of light-matter interactions, leading to an averaging of phonon fluctuations and suppression of phonon-induced dynamic disorder. Here, we show that multilayered materials can extend the exciton-polariton coherence lifetime, up to an order of magnitude, and enhance exciton-polariton transport, which is relevant for developing polariton-based quantum devices [5, 14, 17].
@@
Theory. To explore the dynamics of exciton-polaritons formed in multilayered materials, we consider a light-matter Hamiltonian beyond the long-wavelength approximation [15, 7, 18, 19, 20], which can be derived from the non-relativistic quantum electrodynamics Hamiltonian in the Coulomb gauge [15, 21], written as (in atomic units)
@@
where $\hat{H}_{\text{ex}}$, $\hat{H}_{\text{phn}}$ and $\hat{H}_{\text{cav}}$ are the bare excitonic, phonon, and cavity Hamiltonians, with $\hat{H}_{\mathrm{ex-phn}}$ and $\hat{H}_{\text{ex-cav}}$ describing the exciton-phonon and exciton-cavity interactions. Here, the last term $\hat{H}_{\text{loss}}$ describes cavity photon loss. The bare excitonic Hamiltonian is written as
@@
where $\hat{X}^{\dagger}_{n,m}$ creates an exciton at the site $n$ in the $m$th layer, $\epsilon_{0}$ is the onsite energy and $\tau$ is the hopping parameter. These layers are stacked parallel to each other with an interlayer spacing of $a_{z}=4$ nm, with each site separated laterally by a distance of $a=1.2$ nm, which corresponds to typical perovskite materials [22]. In this work, we also consider one phonon degree of freedom (DOF) per site, with the bare phonon Hamiltonian given by
$\hat{H}_{\text{phn}}=\sum_{n,m}\left(\frac{\hat{p}_{n,m}^{2}}{2}+\frac{1}{2}\omega^{2}\hat{q}_{n,m}^{2}\right),$ (3)
@@
where $\gamma$ is the exciton-phonon coupling strength. The bare cavity Hamiltonian $\hat{H}_{\text{cav}}$ describes a set of confined radiation modes in a Fabry-Pérot optical cavity [26, 7, 27, 28] such that
@@
where $\hat{a}^{\dagger}_{\boldsymbol{k}}$ creates a photon of wavevector ${\boldsymbol{k}}$ with a frequency $\omega_{\boldsymbol{k}}=\frac{c}{\eta}|{\boldsymbol{k}}|$ where $c$ and $\eta=2.4$ are the speed of light and the refractive index, respectively. In this work, we consider only two directions, i.e. $x$ and $y$ such that ${\boldsymbol{k}}=k_{x}\vec{x}+k_{y}\vec{y}$ with $y$ as the cavity quantization direction. Similar to recent work [20, 15, 24, 29, 2, 9, 30] we impose a periodic boundary condition in the $x$ direction, effectively quantizing the $k_{x}=\frac{2\pi n_{y}}{N\cdot a}$ where $n_{x}=0,\pm 1,\pm 2,...$ and consider primary cavity mode along the $y$ direction such that $k_{y}=\frac{\pi}{L}$ with $L=1000~{}\mathrm{\AA}$ as the distance between the two reflective mirrors of the optical cavity. To simplify our notation, we denote $k=k_{x}$ and label all photonic operators and related parameters with $k$, as $k_{y}$ is fixed, such that $\hat{H}_{\text{cav}}=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}\hat{a}^{\dagger}_{\boldsymbol{k}}\hat{a}_{\boldsymbol{k}}\rightarrow\sum_{{k}}\omega_{{k}}\hat{a}^{\dagger}_{{k}}\hat{a}_{{k}}$ and $\omega_{k}=\frac{c}{\eta}\sqrt{k_{z}^{2}+k^{2}}$.
The light-matter interactions beyond the long-wavelength approximation [2, 30, 31, 24] is described by the $\hat{H}_{\text{ex-cav}}$ as
$\hat{H}_{\text{ex-cav}}=\sum_{n,m,{k}}\frac{\Omega_{k}}{\sqrt{N}}\left(\hat{X}_{n,m}^{\dagger}\hat{a}_{k}e^{i{k}x_{n}}+h.c.\right)\sin(k_{y}\cdot y_{m})\,,$ (6)
where $\Omega_{k}=\sqrt{\frac{\omega_{0}}{\omega_{k}}}\Omega_{0}$ ($\Omega_{0}=480$ meV chosen here) is the light-matter coupling strength with the spatial location of the exciton $\hat{X}^{\dagger}_{n,m}$ as $R_{n,m}=x_{n}\vec{x}+y_{m}\vec{y}$. Finally, we model the cavity photon loss using a non-Hermitian Hamiltonian [20, 32, 33, 34]
@@
Figure 2: Exciton-polariton dynamics and spectra in multilayered materials ($M=25$) and the phonon fluctuation synchronization effect. The exciton-polariton angle-resolved spectra at (a) $\gamma=\gamma_{0}$ and (b) $\gamma=2\gamma_{0}$. Exciton-polariton group velocity computed from quantum dynamical simulations vs. theoretical curve with (c) light-matter coupling per layer $\Omega_{k}$ a constant and (d) the total Rabi-splitting $\sqrt{\mathcal{S}}\Omega_{k}$ a constant. Here the phonon coupling $\gamma$ is written in terms of $\gamma_{0}$, a constant. (e) The ratio of group velocities in multilayered material to a single-layer material. (f) Distribution of on-site excitonic energy. Illustration of hopping dynamics in (g) a single-layer material with higher disorder and (h) in a multilayered material with effectively lower disorder. (i) Group velocity in single-layered and multilayered material with constrained sampling of ${q}_{n,m}(t=0)={q}_{n,1}(t=0)$. ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u.)
Quantum Dynamical Approach. In this work, we employ a custom mixed-quantum-classical method based on the mean-field Ehrenfest approach to simulate the dynamics of the light-matter hybrid system. Within the standard Ehrenfest approach, that has been extensively used to simulate non-adiabatic dynamics of polaritons, [35, 36, 7, 37, 38, 9, 20, 10] the nuclear (or slow) degree of freedom is evolved classically, with $\{\hat{q}_{n,m},\hat{p}_{n,m}\}\rightarrow\{{q}_{n,m},{p}_{n,m}\}$, following the Hamiltonian’s equation of motion,
@@
where $|{\Psi}(t)\rangle$ is the electronic-photonic wavefunction at time $t$. The electronic-photonic wavefunction is evolved using the time-dependent Schrödinger equation written as
$i|\dot{\Psi}(t)\rangle$ $=\Bigg{[}\hat{H}_{\mathrm{LM}}-\frac{1}{2}\sum_{n,m}\frac{p_{n,m}^{2}}{2}\Bigg{]}|{\Psi(t)}\rangle.$ (8)
In this work, we confine the exciton-polariton dynamics to the single excited subspace, such that
@@
$\equiv\sum_{k}c_{k}(t)|1_{k}\rangle+\sum_{n,m}{b}_{n,m}(t)|n,m\rangle$ (10)
where $c_{k}(t)$ and ${b}_{n,m}(t)$ are time-dependent coefficients. In the second line, we have introduced the compact representation $|1_{k}\rangle\equiv\hat{a}^{\dagger}_{k}|\bar{0}\rangle$ and $|n,m\rangle\equiv\hat{X}^{\dagger}_{n,m}|\bar{0}\rangle$, with $|\bar{0}\rangle$ as the ground (or vacuum) state of the system, for simplicity. Despite the mixed quantum-classical treatment of the full light-matter Hamiltonian, numerically solving Eq. 8 is extremely challenging as the present work requires a basis of size $\sim 10^{6}$. To resolve this issue, we develop a split-operator approach where a short-time (a single time-step) propagation of $|{\Psi}(t)\rangle$ is obtained as
@@
$=\hat{U}_{\mathrm{ft}}\hat{U}_{\mathrm{B}}\cdot e^{-i(\hat{U}_{\mathrm{B}}^{\dagger}\hat{U}_{\mathrm{ft}}^{\dagger}\hat{H}_{\mathrm{EP}}\hat{U}_{\mathrm{ft}}\hat{U}_{\mathrm{B}})\delta t}\cdot\hat{U}_{\mathrm{B}}^{\dagger}\hat{U}_{\mathrm{ft}}^{\dagger}\cdot e^{-i\hat{H}_{\mathrm{env}}\delta t}|{\Psi}(t)\rangle,$
where $\hat{H}_{\mathrm{LM}}=\hat{H}_{\mathrm{EP}}+\hat{H}_{\mathrm{env}}$ with $\hat{H}_{\mathrm{env}}=\hat{H}_{\text{phn}}+\hat{H}_{\mathrm{ex-phn}}+\hat{H}_{\text{loss}}$ and $\hat{H}_{\mathrm{EP}}=\hat{H}_{\text{ex}}+\hat{H}_{\text{cav}}+\hat{H}_{\text{ex-cav}}$. Here $\hat{H}_{\mathrm{env}}$ is the diagonal in the exciton-photon basis $\{|1_{k}\rangle,|n,m\rangle\}$ chosen here. As a result, the action of the matrix $e^{-i\hat{H}_{\mathrm{env}}\delta t}$ on the vector $|{\Psi}(t)\rangle$ reduces to a simple Hadamard product between a vector containing the diagonal elements of $e^{-i\hat{H}_{\mathrm{env}}\delta t}$ and $|{\Psi}(t)\rangle$. Meanwhile, $\hat{U}_{\mathrm{ft}}$ is a unitary operator that Fourier transforms the excitonic subspace within each layer, such that $\hat{U}_{\mathrm{ft}}|{\Psi}(t)\rangle=\sum_{k}c_{k}(t)|1_{k}\rangle+\hat{U}_{\mathrm{ft}}\sum_{n,m}{b}_{n,m}(t)|n,m\rangle=\sum_{k}c_{k}(t)|1_{k}\rangle+\sum_{k,m}{b}_{k,m}(t)|k,m\rangle$. We perform this Fourier transformation using the Fast Fourier Transformation (FFT) algorithm, which scales as $N\log N$ resulting in a significant reduction of the computation cost [39, 40, 29]. Importantly, here we introduce the unitary operator $\hat{U}_{\mathrm{B}}=\hat{U}_{\mathrm{D}}\cdot\hat{U}_{\mathrm{db}}$, where $\hat{U}_{\mathrm{db}}$ transforms the excitonic subspace into a dark-bright layers subspace [15] and $\hat{U}_{\mathrm{D}}$ diagonalizes this transformed exciton-polariton Hamiltonian,
$\hat{U}_{\mathrm{B}}^{\dagger}\hat{U}_{\mathrm{ft}}^{\dagger}\hat{H}_{\mathrm{EP}}\hat{U}_{\mathrm{ft}}$ $\hat{U}_{\mathrm{B}}=\hat{U}_{\mathrm{B}}^{\dagger}\big{[}\hat{U}_{\mathrm{ft}}^{\dagger}(\hat{H}_{\text{ex}}+\hat{H}_{\text{ex-cav}})\hat{U}_{\mathrm{ft}}+\hat{H}_{\text{cav}}\big{]}\hat{U}_{\mathrm{B}}$ (12)
@@
where $\hat{X}^{\dagger}_{k,b}=\frac{1}{\sqrt{\mathcal{S}}}\sum_{m}\sin(k\cdot y_{m})\hat{X}^{\dagger}_{k,m}$ are bright layer exciton operators, with the normalization constant $\mathcal{S}=\sum_{m}\sin^{2}(k\cdot y_{m})$. Here, $\hat{X}^{\dagger}_{k,b}$ creates an exciton delocalized over all layers with an in-plane wavevector $k$. On the other hand, $\hat{X}^{\dagger}_{k,d}=\sum_{m}s_{m,d}\hat{X}^{\dagger}_{k,m}$, with $\sum_{m}s_{m,d}^{2}=1$ and $\sum_{m}s_{m,d}\cdot\sin(k\cdot y_{m})=0$, are dark layers exciton operators that do not couple to cavity radiation modes. These dark exciton operators form dark exciton bands, which are illustrated in Fig. 1d. In the last line of Eq. 12, we have introduced the upper and lower polariton operators,
$\hat{P}_{k,+}^{\dagger}$ $=\sin\theta_{k}\cdot\hat{a}^{\dagger}_{k}+\cos\theta_{k}\cdot\hat{X}^{\dagger}_{k,b}$ (14)
$\hat{P}_{k,-}^{\dagger}$ $=\cos\theta_{k}\cdot\hat{a}^{\dagger}_{k}-\sin\theta_{k}\cdot\hat{X}^{\dagger}_{k,b}$ (15)
@@
where $\theta_{k}=\frac{1}{2}\tan^{-1}[2\sqrt{\mathcal{S}}\Omega_{k}/(\omega_{k}-\epsilon_{k})]$ is a light-matter mixing angle [41, 7]. This diagonal form allows us to propagate the exciton polariton Hamiltonian in an efficient manner. Finally, an expectation value of an operator $\hat{A}$ is computed as $\langle\hat{A}\rangle\approx\big{\langle}\langle{\Psi}(t)|\hat{A}|{\Psi}(t)\rangle\big{\rangle}_{\mathrm{MFE}}$ where $\langle...\rangle_{\mathrm{MFE}}$ indicates averaging over trajectories generated from the initial realizations of nuclear coordinates $\{{q}_{n,m}(0),{p}_{n,m}(0)\}$. Details of our propagation scheme are provided in the Supplementary Information (SI).

### v2

@@ Abstract
Coupling excitons with quantized radiation has been shown to enable coherent ballistic transport at room temperature inside optical cavities. Previous theoretical works employ a simple description of the material, depicting it as a one-dimensional single layer placed in the middle of an optical cavity, thereby ignoring the spatial variation of the radiation field. In contrast, in most experiments, the optical cavity is filled with organic molecules or multiple layers of two-dimensional materials. Here, we develop an efficient mixed-quantum-classical approach, introducing a bright layer description, to simulate the exciton-polariton quantum dynamics in three dimensions. Our simulations reveal that, for the same Rabi splitting, a multilayered material extends the quantum coherence lifetime and enhances transport compared to a single-layer material. We find that this enhanced coherence can be traced to a synchronization of phonon fluctuations over multiple layers, wherein the collective light-matter coupling in a multilayered material effectively suppresses the phonon-induced dynamical disorder.
@@
However, these theoretical works adopt a simplified light-matter system, where a single layer, modeled as an one-dimensional chain, placed at the center of a two-dimensional optical cavity. This is in contrast to the experiments [2, 3, 6, 5, 15, 16] where the space between the reflective mirrors is filled with molecules or multiple layers of two-dimensional materials, as schematically illustrated in Fig. 1a-b. Consequently, the density of states in these two setups (i.e., multilayered vs. single-layered) is drastically different (see insets in Fig. 1c-d), with a multilayered material featuring an ensemble of optically dark bands that are absent in a single-layered material when coupled to a cavity. Therefore, it is anticipated that the polariton quantum dynamics in these two setups will differ. However, the role of a multilayer configuration in polaritonic quantum dynamics has remained unexplored, which is the focus of this present work.
@@
Figure 1: Schematics and band structures of single-layer vs. multilayer materials. Illustration of (a) a single-layer material and (b) a multilayered material in a three-dimensional optical cavity, respectively. The exciton-polariton band structures of (c) single-layer and (d) multilayered materials in the absence of phonons ($\gamma=0$) in a two-dimensional optical cavity. In a multilayer setup, in addition to the polariton bands, an ensemble of dark (material) bands exists. These dark bands increase the density of states (DOS) near the onsite energy value, which are plotted as insets within (c)-(d).
@@
Simulating exciton-polariton dynamics in multilayered material is an extremely challenging task, as the number of required unit cells in a single layer exceeds $\approx 10^{4}-10^{8}$, thus requiring us to perform a quantum dynamics simulation involving $\approx 10^{6}-10^{10}$ states for $\approx 10^{2}$ layers, typical in these systems. To resolve this fundamental challenge, here we introduce an efficient mixed quantum-classical approach that uses a bright-layer unitary transformation of the light-matter interaction and propagates the dynamics utilizing a split-operator approach. We emphasize, that this approach do not introduce any additional approximation. Using this convenient theoretical tool, we investigate the polaritonic quantum dynamics in multilayered materials. We discover a phonon fluctuation synchronization effect where the exciton-polaritons experience much lower phonon fluctuations and remain coherent for a longer time. We show that this effect originates from the delocalized nature of light-matter interactions, leading to an averaging of phonon fluctuations and suppression of phonon-induced dynamical disorder. Here, we show that multilayered materials can extend the exciton-polariton coherence lifetime, up to an order of magnitude, and enhance exciton-polariton transport, which is relevant for developing polariton-based quantum devices [5, 14, 17].
@@
Theory. To explore the dynamics of exciton-polaritons formed in multilayered materials, we consider a light-matter Hamiltonian beyond the long-wavelength approximation [15, 7, 18, 19, 20]. This Hamiltonian can be derived from the non-relativistic quantum electrodynamics Hamiltonian in the Coulomb gauge [15, 21], written as (in atomic units)
@@
where $\hat{H}_{\text{ex}}$, $\hat{H}_{\text{phn}}$ and $\hat{H}_{\text{cav}}$ are the bare excitonic, phonon, and cavity Hamiltonians, with $\hat{H}_{\mathrm{ex-phn}}$ and $\hat{H}_{\text{ex-cav}}$ describing the exciton-phonon and exciton-cavity interactions. Note that, for the sake of simplicity, we first restrict our discussion to a simple stacked 1D layered system (extending in $x$ and $z$ directions), as we arrive at the same conclusions when considering the full three dimensional problem as shown in Fig. 5. The full three dimensional Hamiltonian is provided in SI. Also note that, as schematically illustrated in Fig. 1, we consider the cavity mirrors to be distributed Bragg reflectors (DBR), so we can ignore the formation of surface plasmon polaritons, which exist at metal-dielectric interfaces [22]. Here, the last term $\hat{H}_{\text{loss}}$ describes cavity photon loss. The bare excitonic Hamiltonian is a one-dimensional tight-binding model with one exciton per unit cell, which is given by
@@
where $\hat{X}^{\dagger}_{n,m}$ creates an exciton at the site $n$ in the $m$th layer, $\epsilon_{0}$ is the onsite energy and $\tau$ is the nearest-neighbor hopping parameter. These layers are stacked parallel to each other with an interlayer spacing of $a_{z}=4$ nm, with each site separated laterally along the $x$ direction by a distance of $a=1.2$ nm, which corresponds to typical perovskite materials [23]. Therefore, the spatial location of the exciton $\hat{X}^{\dagger}_{n,m}$ is $R_{n,m}=x_{n}\vec{e}_{x}+z_{m}\vec{e}_{z}=na\vec{e}_{x}+ma_{z}\vec{e}_{z}$. In this work, we also consider one phonon degree of freedom per site, with the bare phonon Hamiltonian given by
$\hat{H}_{\text{phn}}=\sum_{n,m}\frac{1}{2}\left({\hat{p}_{n,m}^{2}}+\omega^{2}\hat{q}_{n,m}^{2}\right),$ (3)
@@
where $\gamma$ is the exciton-phonon coupling strength. The bare cavity Hamiltonian $\hat{H}_{\text{cav}}$ describes a set of confined radiation modes, considering only transverse magnetic (TM) polarized mode (we consider both polarizations in the full three dimensional scenario in Fig. 5), in a Fabry-Pérot optical cavity [27, 7, 28, 29] such that
@@
where $\hat{a}^{\dagger}_{\boldsymbol{k}}$ creates a photon of wavevector $\boldsymbol{k}$} with a frequency $\omega_{\boldsymbol{k}}=\frac{c}{\eta}|{\boldsymbol{k}}|$ where $c$ and $\eta=2.4$ are the speed of light and the refractive index, respectively. As noted earlier, in the simplified scenario, we consider two directions $x$ (longitudinal) and $z$ (transverse), such that ${\boldsymbol{k}}=k_{x}\vec{e}_{x}+k_{z}\vec{e}_{z}$ with $z$ as the cavity quantization direction. Similar to recent work [20, 15, 25, 30, 2, 9, 31] we impose a periodic boundary condition in the $x$ direction, effectively quantizing the $k_{x}=\frac{2\pi n_{x}}{Na}$ where $n_{x}=0,\pm 1,\pm 2,$.... Here, we consider first cavity quantization mode along the $z$ direction such that $k_{z}=\frac{\pi}{L}$ with $L=100$ nm as the distance between the two reflective mirrors of the optical cavity. To simplify our notation we denote $k=k_{x}$ and label all photonic operators and related parameters with $k$, as $k_{z}$ is fixed, such that $\hat{H}_{\text{cav}}=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}\hat{a}^{\dagger}_{\boldsymbol{k}}\hat{a}_{\boldsymbol{k}}\rightarrow\sum_{{k}}\omega_{{k}}\hat{a}^{\dagger}_{{k}}\hat{a}_{{k}}$ and $\omega_{k}=\frac{c}{\eta}\sqrt{k_{z}^{2}+k^{2}}$.
The light-matter interactions beyond the long-wavelength approximation in the simplified two-dimensional scenario [2, 31, 32, 25] is described by
$\hat{H}_{\text{ex-cav}}=\sum_{n,m,{k}}\frac{\Omega_{k}}{\sqrt{N}}\left(\hat{X}_{n,m}^{\dagger}\hat{a}_{k}e^{i{k}x_{n}}+h.c.\right)\sin(k_{z}z_{m})\,$, (6)
where $\Omega_{k}=\sqrt{\frac{\omega_{0}}{\omega_{k}}}\Omega_{0}$ is the light-matter coupling strength. Throughout the manuscript we use $\Omega_{0}=330$ meV, therefore the light-matter coupling remains within the strong coupling regime and away from the ultra-strong coupling regime. Other data for larger and smaller $\Omega_{0}$ can be found in the Supplementary Information (SI). Finally, we model the cavity photon loss using a non-Hermitian Hamiltonian [20, 33, 34, 35] given by
@@
Figure 2: Exciton-polariton dynamics and spectra in multilayered material ($M=25$) and the phonon fluctuation synchronization effect in a two-dimensional cavity. The exciton-polariton angle-resolved spectra at (a) $\gamma=\gamma_{0}$ and (b) $\gamma=2\gamma_{0}$. Exciton-polariton group velocity computed from quantum dynamical simulations vs. theoretical curve with (c) light-matter coupling per layer $\Omega_{k}$ a constant and (d) the total Rabi-splitting $\sqrt{\mathcal{S}}\Omega_{k}$ a constant. Here, the phonon coupling $\gamma$ is written in terms of $\gamma_{0}$, a constant. (e) The ratio of group velocities in a multilayered material to a single-layer material. (f) Distribution of on-site excitonic energy. Illustration of hopping dynamics in (g) a single-layer material with higher disorder and (h) in a multilayered material with effectively lower disorder. (i) Group velocity in single-layered and multilayered material with constrained sampling of ${q}_{n,m}(t=0)={q}_{n,1}(t=0)$. ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u., $\Omega_{0}=330$ meV, and 60000 sites.)
Figure 3: Comparing exciton-polariton dynamics in single-layer vs. multilayered materials ($M=25$) in a two-dimensional cavity with the initial excitation at $E_{0}=2.62$ eV. (a) Time-dependent exciton-polariton population for both single-layer and multilayered material with no exciton-phonon couplings ($\gamma=0$). (b)-(c) Same as (a) but in the presence of phonon couplings $\gamma=2\gamma_{0}$ in (b) a single and (c) multilayered material. (d) Site-resolved exciton density at $t=0.75$ ps and $\gamma=2\gamma_{0}$. To determine the multilayer exciton populations in (d), we summed the density contributions from all layers at each site. (e)-(h) Site and layer-resolved exciton density for (e)-(f) a single layer and (g)-(h) multilayered material at $t=0$ and $t=0.75$ ps with the same parameters as in (d). ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u., $\Omega_{0}=330$ meV, and 60001 sites.)
Quantum Dynamical Approach. In this work, we employ a custom mixed-quantum-classical method based on the mean-field Ehrenfest approach to simulate the dynamics of the light-matter hybrid system. Within the standard Ehrenfest approach, that has been extensively used to simulate non-adiabatic dynamics of polaritons, [36, 37, 7, 38, 39, 9, 20, 10, 40, 41] the phonon (or slow) degree of freedom is evolved classically, with $\{\hat{q}_{n,m},\hat{p}_{n,m}\}\rightarrow\{{q}_{n,m},{p}_{n,m}\}$, following the Hamiltonian’s equation of motion,
@@
where $|{\Psi}(t)\rangle$ is the excitonic-photonic wavefunction at time $t$. The excitonic-photonic wavefunction is evolved using the time-dependent Schrödinger equation written as
$i|\dot{\Psi}(t)\rangle$ $=\Bigg{[}\hat{H}_{\mathrm{LM}}-\sum_{n,m}\frac{p_{n,m}^{2}}{2}\Bigg{]}|{\Psi(t)}\rangle.$ (8)
In this work, we confine the exciton-polariton dynamics to the single excited subspace. Therefore, the exciton-photon wavefunction is given
@@
$=\sum_{k}c_{k}(t)|1_{k}\rangle+\sum_{n,m}{b}_{n,m}(t)|n,m\rangle$ (10)
where $c_{k}(t)$ and ${b}_{n,m}(t)$ are time-dependent coefficients. In the second line, we have introduced the compact representation $|1_{k}\rangle\equiv\hat{a}^{\dagger}_{k}|\bar{0}\rangle$ and $|n,m\rangle\equiv\hat{X}^{\dagger}_{n,m}|\bar{0}\rangle$, with $|\bar{0}\rangle$ as the ground (or vacuum) state of the system, for simplicity. Despite the mixed quantum-classical treatment of the full light-matter Hamiltonian, numerically solving Eq. 8 is extremely challenging as the present work requires a basis of size $\approx 10^{6}$ for the simplified two-dimensional scenario ($\approx 10^{8}$ presented in Fig. 5). To resolve this issue, we develop a split-operator approach where a short-time (a single time-step) propagation of $|{\Psi}(t)\rangle$ is obtained as
@@
$=\hat{U}_{\mathrm{ft}}\hat{U}_{\mathrm{B}}e^{-i(\hat{U}_{\mathrm{B}}^{\dagger}\hat{U}_{\mathrm{ft}}^{\dagger}\hat{H}_{\mathrm{EP}}\hat{U}_{\mathrm{ft}}\hat{U}_{\mathrm{B}})\delta t}\hat{U}_{\mathrm{B}}^{\dagger}\hat{U}_{\mathrm{ft}}^{\dagger}e^{-i\hat{H}_{\mathrm{env}}\delta t}|{\Psi}(t)\rangle,$
where $\hat{H}_{\mathrm{LM}}=\hat{H}_{\mathrm{EP}}+\hat{H}_{\mathrm{env}}$ with $\hat{H}_{\mathrm{env}}=\hat{H}_{\text{phn}}+\hat{H}_{\mathrm{ex-phn}}+\hat{H}_{\text{loss}}$ and $\hat{H}_{\mathrm{EP}}=\hat{H}_{\text{ex}}+\hat{H}_{\text{cav}}+\hat{H}_{\text{ex-cav}}$. Here $\hat{H}_{\mathrm{env}}$ is the diagonal in the exciton-photon basis $\{|1_{k}\rangle,|n,m\rangle\}$ chosen here. As a result, the action of the matrix $e^{-i\hat{H}_{\mathrm{env}}\delta t}$ on the vector $|{\Psi}(t)\rangle$ reduces to a simple Hadamard product between a vector containing the diagonal elements of $e^{-i\hat{H}_{\mathrm{env}}\delta t}$ and $|{\Psi}(t)\rangle$. Meanwhile, $\hat{U}_{\mathrm{ft}}^{\dagger}$ is a unitary operator that Fourier transforms the excitonic subspace within each layer, such that $\hat{U}^{\dagger}_{\mathrm{ft}}|{\Psi}(t)\rangle=\sum_{k}c_{k}(t)|1_{k}\rangle+\hat{U}^{\dagger}_{\mathrm{ft}}\sum_{n,m}{b}_{n,m}(t)|n,m\rangle=\sum_{k}c_{k}(t)|1_{k}\rangle+\sum_{k,m}{b}_{k,m}(t)|k,m\rangle$. We perform this Fourier transformation using the Fast Fourier Transformation (FFT) algorithm, which scales as $N\log N$, resulting in a significant reduction of the computation cost [42, 43, 30]. Importantly, here we introduce the unitary operator $\hat{U}_{\mathrm{B}}=\hat{U}_{\mathrm{db}}\hat{U}_{\mathrm{D}}$, where $\hat{U}_{\mathrm{db}}$ transforms the excitonic subspace into a dark-bright layers subspace [15] and $\hat{U}_{\mathrm{D}}$ diagonalizes this transformed exciton-polariton Hamiltonian. This unitary transformation is given by (a much more involved unitary transformation in case of the full 3D scenario is provided in the SI)
$\hat{U}_{\mathrm{B}}^{\dagger}\hat{U}_{\mathrm{ft}}^{\dagger}\hat{H}_{\mathrm{EP}}\hat{U}_{\mathrm{ft}}$ $\hat{U}_{\mathrm{B}}=\hat{U}_{\mathrm{B}}^{\dagger}\big{[}\hat{U}_{\mathrm{ft}}^{\dagger}(\hat{H}_{\text{ex}}+\hat{H}_{\text{ex-cav}})\hat{U}_{\mathrm{ft}}+\hat{H}_{\text{cav}}\big{]}\hat{U}_{\mathrm{B}}$
@@
where $\hat{X}^{\dagger}_{k,b}=\frac{1}{\sqrt{\mathcal{S}}}\sum_{m}\sin(k_{z}z_{m})\hat{X}^{\dagger}_{k,m}$ are bright layer exciton operators, with the normalization constant $\mathcal{S}=\sum_{m}\sin^{2}(k_{z}z_{m})$. Here, $\hat{X}^{\dagger}_{k,b}$ creates an exciton delocalized over all layers with an in-plane wavevector $k$. On the other hand, $\hat{X}^{\dagger}_{k,d}=\sum_{m}s_{m,d}\hat{X}^{\dagger}_{k,m}$ is a dark layers exciton operator, with $\sum_{m}s^{*}_{m,d}s_{m,d^{\prime}}=\delta_{d,d^{\prime}}$ ($\delta_{d,d^{\prime}}$ is the Kronecker delta) and $\frac{1}{\sqrt{\mathcal{S}}}\sum_{m}s_{m,d}\sin(k_{z}z_{m})=0$. Mathematically, in the absence of phonons, these dark layers’ exciton operators do not couple to cavity radiation modes, and form dark exciton bands, which are illustrated in Fig. 1d. In the last line of Eq. Exciton-Polariton Dynamics in Multilayered Materials, we have introduced the upper and lower polariton operators,
$\hat{P}_{k,+}^{\dagger}$ $=\sin\theta_{k}\hat{a}^{\dagger}_{k}+\cos\theta_{k}\hat{X}^{\dagger}_{k,b}$ (13)
$\hat{P}_{k,-}^{\dagger}$ $=\cos\theta_{k}\hat{a}^{\dagger}_{k}-\sin\theta_{k}\hat{X}^{\dagger}_{k,b}$ (14)
@@
where $\theta_{k}=\frac{1}{2}\tan^{-1}[2\sqrt{\mathcal{S}}\Omega_{k}/(\omega_{k}-\epsilon_{k})]$ is a light-matter mixing angle [44, 7]. This diagonal form enables us to propagate the exciton polariton Hamiltonian in an efficient manner. Finally, expectation value of an operator $\hat{A}$ is computed as $\langle\hat{A}\rangle\approx\big{\langle}\langle{\Psi}(t)|\hat{A}|{\Psi}(t)\rangle\big{\rangle}_{\mathrm{MFE}}$ where $\langle...\rangle_{\mathrm{MFE}}$ indicates averaging over trajectories generated from the initial realizations of nuclear coordinates $\{{q}_{n,m}(0),{p}_{n,m}(0)\}$. Details of our propagation scheme are provided in the SI.